OpenStudy (anonymous):

Determine which ordered pairs are solution for y = 5x - 4. A. (2,-6) B. (0,-4) C. (0,-5) D. (4/5,0)

Directrix (directrix):

At first inspection, it appears that B. (0,-4) and D. (4/5,0) are both solutions. Let's check:

OpenStudy (anonymous):

i think your right

Directrix (directrix):

y = 5x - 4 -4 = 5(0) -4 so -4 = -4 Yes, a solution 0 = 5(4/5) -4 ?? 0 = 20/5 - 4 = 4-4 = 0 Yes, a solution

OpenStudy (anonymous):

so its B and D, Correct?

Directrix (directrix):

Yes.

OpenStudy (anonymous):

ok